5^log(base10)x=5 + 4x^log(base10)5. find the value of x?




Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
let log(x)=y
=>5^y=5+4*5^y
=>5^y=-5/3
can't be negative
can you clearly state the question